OpenStudy (anonymous):

A car travels 172 miles using 7 gallons of gas. At that rate, how far can the car travel using 42 gallons of gas?

OpenStudy (whpalmer4):

You could do this as a proportion problem, or as a rate problem. Proportion: \[\frac{172 \text{ miles}}{7 \text{ gallons}} = \frac{x \text{ miles}}{42 \text{ gallons}}\]Cross-multiply to get \[172*42 = 7*x\]\[x = \text{ ?}\] As a rate: \[172 \text{ miles} / 7 \text{ gallons} \approx 24.6 \text { miles/gallon}\] Distance = rate * fuel \[D = 24.6 \text{ miles/gallon} * 42 \text{ gallon} = \]